# Less is More: Compact Clue Selection for Efficient Retrieval-Augmented Generation Reasoning

### v6

@@ 4.2.2. Inference
Given the query $q$ and the extracted sentences $\mathcal{S}^{c}$, the $Reranker$ model calculates the relevance score between every sentence $s_{j}^{c}\in\mathcal{S}^{c}$ and query $q$. The reranked clues are defined as:
@@ 4.3. Adaptive Truncator
The adaptive truncator aims to capture the minimal necessary clues based on the complexity of the question and the documents, ensuring sufficient clues for accurate answer generation.
@@ 4.3.1. Training
To determine the optimal clues subset $\mathcal{S}^{t}$ for each query $q$, we perform data annotation based on the reranked answer clues $\mathcal{S}^{r}$ obtained from the previous reranking step. Given a query $q$ and its reranked clues $\mathcal{S}^{r}=\{s_{1}^{r},\dots,s_{n}^{r}\}$, the objective is to identify the smallest subset $\mathcal{S}^{t}$ such that the RAG system’s generation model $M$ can generate the correct answer $y$ based on $q$ and $\mathcal{S}^{t}$. We define $D_{k}=\{s_{1}^{r},\dots,s_{k}^{r}\}$, where $1\leq k\leq n$. The performance on each subset $D_{k}$ is evaluated by checking if the generation model’s output $M(q,D_{k})$ matches the ground truth $y$. The correctness condition is defined as:
@@
During the model training stage, we fine-tune an LLM based on the data annotations. The $Truncator$ is trained to predict the minimal subset of reranked clues $S_{t}\subseteq S_{r}$ that suffices to answer the query:
@@ 4.3.2. Inference
During inference, given a new query $q$ and its reranked sentences $\mathcal{S}^{r}$, the $Truncator$ truncates $\mathcal{S}^{r}$ to the minimal sufficient clue set $\mathcal{S}^{t}=\{S^{r}_{1},\dots,S^{r}_{K_{g}}\}$, where $K_{g}$ represents the number of sentences needed to answer the query. The truncation process uses the prompt provided in Appendix A.2. Finally, the RAG generation module concatenates the query $q$ with the filtered answer clues $\mathcal{S}^{t}$ using the following prompt to reason the answer:
<system> You are a helpful, respectful, and honest assistant. Answer the question with couple of words using the provided documents. For example: Question: What is the capital of France? Output: Paris. </system>
@@ 5. Experiments
Table 2. Experimental results on three benchmark datasets using LLaMA3-8B-Instruct and Qwen3-14B as generators. The retriever is DPR. All baselines and our method are conducted using the same test sets and retrieval corpus. SubEM=substring Exact Match, F1 = F1 score, CR = $\frac{\#\text{tokens in retrieved documents}}{\#\text{tokens in compressed text}}$. The best results are in bold and the second are underlined.
@@ 5.1.1. Datasets
We evaluate our method on three QA benchmark datasets, including (1) Open-Domain QA, represented by NaturalQuestions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (TQA) (Joshi et al., 2017); (2) Multi-Hop QA, represented by HotpotQA (HQA) (Yang et al., 2018). Table 11 in Appendix B.1 illustrates the statistics of these datasets.
@@ 5.1.2. Evaluation Metrics
Since answer style mismatch may cause additional variance, we follow prior work (Zhu et al., 2024b; Peng et al., 2025; Xu et al., 2024a) and adopt Substring Exact Match (SubEM) and F1 for evaluation. SubEM checks whether the gold answer appears as a substring in the prediction, while F1 measures token-level overlap with the reference. For efficiency, we report the compression ratio (CR), defined as the ratio of original to compressed context length. We also measure Total Latency, including offline preprocessing and online answer generation, and Online Latency, the time from user query submission to answer generation using the preprocessed results, which can provide a more comprehensive assessment of practical efficiency.
@@ 5.1.3. Baselines
Reranking Methods: (i) BM25 (Robertson et al., 2009), a classic lexical matching method that scores and ranks documents based on term frequency, inverse document frequency, and document length normalization; (ii) BGE-reranker (Xiao et al., 2024), a neural reranker that computes dense embeddings for queries and documents, ranking documents according to semantic similarity in the embedding space; (iii) RECOMP-extr (Xu et al., 2024b), which employs a fine-tuned cross-encoder to select salient sentences through dense retrieval.
@@
Abstractive Methods: (i) RECOMP-abs (Xu et al., 2024b), which uses a T5-based model to perform abstractive summarization, compressing documents into shorter token sequences through autoregressive generation; (ii) Refiner (Li et al., 2024), which leverages large LLMs to extract and structure query-relevant content from retrieved documents, producing a hierarchical output based on intrinsic document knowledge; (iii) BottleNeck (Zhu et al., 2024a), which employs reinforcement learning and information bottleneck theory to improve both filtering and generation.
@@ 5.1.4. Implementation Details
We use LLaMA3-8B-Instruct  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Dubey et al., 2024), Qwen3-14B  [^4] https://huggingface.co/Qwen/Qwen3-14B (Yang et al., 2025), and LLaMA3.3-70B-Instruct  [^5] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct (Dubey et al., 2024) (presented in Appendix C) as the generators, covering medium, large, and ultra-large LLMs. To ensure high coverage and quality of retrieved information (Cuconasu et al., 2024), we follow prior work (Zhu et al., 2024a; Xu et al., 2024b; Li et al., 2024; Wang et al., 2023; Zhao et al., 2024) and utilize the adversarial Dense Passage Retriever (DPR) (Karpukhin et al., 2020) to retrieve Top-5 passages from the full Wikipedia passages for each query. All baselines and our method are conducted using the same test sets and retrieval corpus, which guarantees consistency between baseline methods and our approach. To further reduce computational cost and latency, our clue extractor and clue truncator are based on LLaMA3.2-3B-Instruct  [^6] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct (Dubey et al., 2024) and we train the model using the LoRA method (Hu et al., 2021) within the LLaMAFactory framework (Zheng et al., 2024). For clue reranker, we implement Sentence-BERT (Reimers and Gurevych, 2020) using distilbert-base-uncased  [^7] https://huggingface.co/distilbert/distilbert-base-uncased. More details are provided in Appendix B.2.
@@ 5.3. Ablation Study
Table 3. Ablation study on NQ, TriviaQA and HotpotQA test set. We use Llama3-8B-Instruct as the generator and SubEM for evaluation.
@@
We analyze the effect of fine-tuning on Clue Extraction by comparing two approaches: 1) Direct Extraction. Using the extractor without fine-tuning to extract clue sentences from retrieved documents based on the given prompt (see Appendix A.1); 2) Fine-tuned Extraction. Using fine-tuned extractor to select answer-containing sentences. As shown in Table 4, Fine-tuned Extraction consistently outperforms Direct Extraction by leveraging task-specific knowledge to identify more relevant sentences.
Table 4. Performance comparison between Direct Extraction and Fine-tuned Extraction across three datasets. We use Llama3-8B-Instruct as the generator.
@@ 5.3.3. Threshold of KNN-Based Extraction
To validate the effectiveness of our adaptive truncation strategy, we compare it with random truncation. As shown in Table 5, our adaptive truncation method outperforms random truncation in both subEM and F1. This improvement occurs because adaptive truncation dynamically selects context based on feedback from the downstream generator, retaining the most relevant and informative content to improve model performance.
Table 5. Performance comparison between Random Truncation and Adaptive Truncation across three datasets. We use Llama3-8B-Instruct as the generator.
@@ 5.4. System Latency Evaluation
To evaluate the overall efficiency of each method, we measured the average total latency (including offline preprocessing and online answer generation) and online latency (the time to answer generation using the preprocessed results) for processing a single sample on the NQ test set, along with the corresponding SubEM scores. This measurement setup provides a more comprehensive assessment of practical efficiency. As shown in Figure 5, RECOMP-extr exhibits the lowest total latency but relatively lower SubEM (40.71%), whereas our method achieves the highest SubEM (46.98%) while maintaining a relatively low total latency (2.14 s, total) and online latency (0.21 s). This indicates that with lightweight reranker and truncator design, our method can improve accuracy without significantly increasing latency, thereby achieving a better trade-off between efficiency and accuracy.
Figure 5. Latency comparison across baselines and our method. The top shows Total Latency along with SubEM performance, while the bottom shows Online Latency. Experiments were conducted on two NVIDIA RTX PRO 6000 GPUs.
@@ 5.5.1. Cascading Errors Resilience Analysis
5.5.2. Performance under Unreliable Retrieval
The truncator not only shortens context but also helps filter out unreliable contents. We conduct experiments on samples from NQ and HotpotQA test set without gold-standard answers to simulate unreliable retrieval. As shown in Table 7, CompSelect’s truncator effectively suppresses noise and prevents answer degradation. This is because CompSelect is trained to allow empty outputs, whereas baseline models perform poorly under unreliable retrieval conditions, as they must choose an output answer clue.
Table 7. Results under unreliable retrieval. We use Llama3-8B-Instruct as the generator and F1 for evaluation. The best results are in bold and the second are underlined.
@@ 5.5.2. Performance under Unreliable Retrieval
Ours (w/ Truncator) 15.36 16.87
@@ 5.6. Cross-Task Generalization
To further evaluate the generalization capability of CompSelect, we conduct experiments on 1,200 randomly sampled instances from a Conversational Multi-Doc QA dataset  [^8] https://sites.google.com/view/wsdm24-docqa, with the model trained solely on HotpotQA. This setup allows us to examine whether the model can adapt to a new conversational QA scenario without direct exposure during training. As shown in Table 8, CompSelect consistently achieves strong performance in this unseen dataset, highlighting its robustness and demonstrating promising generalization across diverse QA tasks and domains.
@@
Ours 0.4123 0.1798 0.2512
6. Conclusion and Future Work
In this work, we propose CompSelect, a compact clue selection mechanism for LLM-centric RAG that efficiently extracts, organizes, and truncates answer-relevant information from large-scale documents.By framing retrieval as a MinMax optimization, its three modules, the clue extractor, the re-ranker, and the truncator, work together to enhance reasoning quality and efficiency. Experiments on three QA datasets show that CompSelect improves performance and reduces latency, while remaining robust to unreliable retrieval and generalizing well across scenarios. Future work could explore integrating it with generative retrieval, allowing models to generate document indices, thereby avoiding costly online processing and further reducing end-to-end latency.
@@ A.1. Prompt for the Clue Extractor
We present our prompt for clue extraction in Table 9, which plays a crucial role in identifying and selecting relevant information from the input documents. This prompt is designed to guide the model in extracting the most informative sentences, those most likely to contain the answer to the given question.
@@ A.2. Prompt for the Adaptive Truncator
We present our prompt for adaptive truncation in Table 10. The prompt is designed to guide the model in optimizing context truncation based on the complexity of the question and the quality of the retrieved documents thereby improving the efficiency of the language model. Specifically, given a question and a ranked list of sentences, the model’s task is to identify and retain the most relevant sentences while truncating those that are irrelevant to the question. Through this process, the model is able to maintain answer accuracy while reducing unnecessary information, thus enhancing processing efficiency.
@@ B.2. More Implementation Details
We use LLaMA3-8B-Instruct (Dubey et al., 2024), Qwen3-14B(Yang et al., 2025), and LLaMA3.3-70B-Instruct (Dubey et al., 2024) as the generators, covering medium, large, and ultra-large LLMs. All of these models demonstrate strong performance across various tasks and exhibit high flexibility. To reduce computational cost and latency, our clue extractor and adaptive truncator are based on LLaMA3.2-3B-Instruct (Dubey et al., 2024). We apply the LoRA method (Hu et al., 2021) for fine-tuning, an efficient low-rank adaptation technique that reduces the computational cost of parameter updates while maintaining model performance. LoRA is applied to both the clue extractor and adaptive truncator. Training is conducted on two NVIDIA RTX PRO 6000 GPUs for 12 epochs. The initial learning rate is set to $1\times 10^{-4}$, and the batch size is 4. Gradient accumulation is employed to simulate larger effective batch sizes and improve training stability. The best model is selected based on validation set performance. During fine-tuning, the models are trained on the three QA datasets using a KNN-based sentence selection method. Data preprocessing is accelerated with 16 parallel workers per training epoch. The maximum input length is set to 4096 tokens to accommodate long-context information. For the clue reranker, we employ Sentence-BERT (Reimers and Gurevych, 2020) with the distilbert-base-uncased model to generate high-quality sentence embeddings for computing sentence similarity. Training uses the Adam optimizer with a batch size of 64, a learning rate of $2\times 10^{-5}$, and 1000 warm-up steps, and runs for 4 epochs.

### v7

@@ 4.2.2. Inference
Given the query $q$ and the extracted sentences $\mathcal{S}^{c}$, the $Reranker$ model calculates the relevance score between every sentence $s_{j}^{c}\in\mathcal{S}^{c}$ and query $q$. The reranked clue set is defined as:
@@ 4.3. Adaptive Truncator
The adaptive truncator is designed to identify the minimal necessary clues based on the complexity of the question and the documents, ensuring sufficient clues for accurate answer generation.
@@ 4.3.1. Training
To determine the optimal clue subset $\mathcal{S}^{t}$ for each query $q$, we perform data annotation based on the reranked answer clues $\mathcal{S}^{r}$ obtained from the previous reranking step. Given a query $q$ and its reranked clues $\mathcal{S}^{r}=\{s_{1}^{r},\dots,s_{n}^{r}\}$, the objective is to identify the smallest subset $\mathcal{S}^{t}$ such that the RAG system’s generation model $M$ can generate the correct answer $y$ based on $q$ and $\mathcal{S}^{t}$. We define $D_{k}=\{s_{1}^{r},\dots,s_{k}^{r}\}$, where $1\leq k\leq n$. The performance on each subset $D_{k}$ is evaluated by checking if the generation model’s output $M(q,D_{k})$ matches the ground truth $y$. The correctness condition is defined as:
@@
During the model training stage, we fine-tune an LLM based on the data annotations. The $Truncator$ is trained to predict the minimal subset of reranked clues $\mathcal{S}^{t}\subseteq\mathcal{S}^{r}$ that suffices to answer the query:
@@ 4.3.2. Inference
During inference, given a new query $q$ and its reranked sentences $\mathcal{S}^{r}$, the $Truncator$ truncates $\mathcal{S}^{r}$ to the minimal sufficient clue set $\mathcal{S}^{t}=\{s^{r}_{1},\ldots,s^{r}_{K_{g}}\}$, where $K_{g}$ represents the number of sentences needed to answer the query. The truncation process uses the prompt provided in Appendix A.2. Finally, the RAG generation module concatenates the query $q$ with the filtered answer clues $\mathcal{S}^{t}$ using the following prompt to reason the answer:
<system> You are a helpful, respectful, and honest assistant. Answer the question with a couple of words using the provided documents. For example: Question: What is the capital of France? Output: Paris. </system>
@@ 5. Experiments
Table 2. Experimental results (%) on three datasets using LLaMA3-8B-Instruct and Qwen3-14B as the generators. The retriever is DPR. All baselines and our method are conducted using the same test sets and retrieval corpus.
@@ 5.1.1. Datasets
We evaluate our method on three QA datasets, including (1) Open-Domain QA, represented by NaturalQuestions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017); (2) Multi-Hop QA, represented by HotpotQA (Yang et al., 2018). Table 11 in Appendix B.1 illustrates the statistics of these datasets.
@@ 5.1.2. Evaluation Metrics
Since answer style mismatch may cause additional variance, we follow prior work (Zhu et al., 2024a; Xu et al., 2024b; Wang et al., 2025b) and adopt Substring Exact Match (SubEM) and F1 for evaluation. SubEM checks whether the gold answer appears as a substring in the prediction, while F1 measures token-level overlap with the reference. For efficiency, we report the compression ratio (CR), defined as the ratio of original to compressed context length. We also measure Total Latency, including offline preprocessing and online answer generation, and Online Latency, the time from user query submission to answer generation using the preprocessed results, which can provide a more comprehensive assessment of practical efficiency.
@@ 5.1.3. Baselines
Reranking Methods: (i) BM25 (Robertson et al., 2009), a classic lexical matching method that scores and ranks documents based on term frequency, inverse document frequency, and document length normalization; (ii) BGE-reranker (Xiao et al., 2024), a neural reranker that computes dense embeddings for queries and documents, ranking documents according to semantic similarity in the embedding space; (iii) RECOMP-extr (Xu et al., 2024a), which employs a fine-tuned cross-encoder to select salient sentences through dense retrieval.
@@
Abstractive Methods: (i) RECOMP-abs (Xu et al., 2024a), which uses a T5-based model to perform abstractive summarization, compressing documents into shorter token sequences through autoregressive generation; (ii) Refiner (Li et al., 2024), which leverages large LLMs to extract and structure query-relevant content from retrieved documents, producing a hierarchical output based on intrinsic document knowledge; (iii) BottleNeck (Zhu et al., 2024b), which employs reinforcement learning and information bottleneck theory to improve both filtering and generation.
@@ 5.1.4. Implementation Details
We use LLaMA3-8B-Instruct  (Dubey et al., 2024), Qwen3-14B (Yang et al., 2025), and LLaMA3.3-70B-Instruct (Dubey et al., 2024) (presented in Appendix C) as the generators, covering medium, large, and ultra-large LLMs. To ensure high coverage and quality of retrieved information (Cuconasu et al., 2024), we follow prior work (Zhu et al., 2024b; Xu et al., 2024a; Li et al., 2024; Zhang et al., 2026) and utilize the adversarial Dense Passage Retriever (DPR) (Karpukhin et al., 2020) to retrieve Top-5 passages from the full Wikipedia passages for each query. All baselines and our method are conducted using the same test sets and retrieval corpus, which guarantees consistency between baseline methods and our approach. To further reduce computational cost and latency, our clue extractor and clue truncator are based on LLaMA3.2-3B-Instruct (Dubey et al., 2024) and we train the model using the LoRA method (Hu et al., 2021) within the LLaMAFactory framework (Zheng et al., 2024). For the clue reranker, we implement Sentence-BERT (Reimers and Gurevych, 2020) using distilbert-base-uncased. More details are provided in Appendix B.2.
@@ 5.3. Ablation Study
Table 3. Ablation study on NQ, TriviaQA and HotpotQA test set. We use LLaMA3-8B-Instruct as the generator and SubEM (%) for evaluation.
@@
We analyze the effect of fine-tuning on clue extraction by comparing two approaches: 1) Direct Extraction, which uses the extractor without fine-tuning to extract clue sentences from retrieved documents based on the given prompt (see Appendix A.1); 2) Fine-tuned Extraction, which uses fine-tuned extractor to select answer-containing sentences. As shown in Table 4, Fine-tuned Extraction consistently outperforms Direct Extraction by leveraging task-specific knowledge to identify more relevant sentences.
Table 4. Performance (%) comparison between Direct Extraction and Fine-tuned Extraction across three datasets. We use LLaMA3-8B-Instruct as the generator.
@@ 5.3.3. Threshold of KNN-Based Extraction
To validate the effectiveness of our adaptive truncation strategy, we compare it with random truncation. As shown in Table 5, our adaptive truncation consistently outperforms random truncation in both SubEM and F1 scores. This improvement arises because adaptive truncation dynamically selects the most relevant context based on real-time feedback from the downstream generator, retaining sentences that are both highly informative and directly pertinent to the query. By doing so, it preserves critical information while reducing unnecessary content, enhancing the model’s answer accuracy and overall reasoning capability. In contrast, random truncation does not account for the importance of individual sentences and may discard key clues, leading to degraded performance.
Table 5. Performance (%) comparison between Random Truncation and Adaptive Truncation across three datasets. We use Llama3-8B-Instruct as the generator. We report SubEM and F1 for evaluation.
@@
Figure 4. Latency comparison across baselines and our method using LLaMA3-8B-Instruct. The top shows Total Latency along with SubEM performance, while the bottom shows Online Latency. Experiments were conducted on two NVIDIA RTX PRO 6000 GPUs.
@@ 5.4. System Latency Evaluation
To evaluate the overall efficiency of each method, we measured the average total latency (including offline preprocessing and online answer generation) and online latency (the time to answer generation using the preprocessed results) for processing a single sample on the NQ test set, along with the corresponding SubEM scores. This measurement setup provides a more comprehensive assessment of practical efficiency. As shown in Figure 4, RECOMP-extr exhibits the lowest total latency but relatively lower SubEM, whereas our method achieves the highest SubEM while maintaining a relatively low total latency and online latency. This indicates that with lightweight reranker and truncator design, our method can improve accuracy without significantly increasing latency, thereby achieving a better trade-off between efficiency and accuracy.
@@ 5.5.1. Cascading Errors Resilience Analysis
Table 7. Results under unreliable retrieval. We use LLaMA3-8B-Instruct as the generator and F1 (%) for evaluation. The best results are in bold and the second are underlined.
@@
[…]
5.5.2. Performance under Unreliable Retrieval
The truncator not only shortens context but also helps filter out unreliable content. We conduct experiments on samples from the NQ and HotpotQA test sets without gold-standard answers to simulate unreliable retrieval. As shown in Table 5.5.1, CompSelect’s truncator effectively suppresses noise and prevents answer degradation. This is because CompSelect is trained to allow empty outputs, whereas baseline models perform poorly under unreliable retrieval conditions, as they must choose an output answer clue.
@@ 5.6. Cross-Task Generalization
To further evaluate the generalization capability of CompSelect, we perform inference on 1,200 randomly sampled instances from a Conversational Multi-Doc QA dataset  [^3] https://sites.google.com/view/wsdm24-docqa, with the model trained solely on HotpotQA dataset. This setup allows us to examine whether the model can adapt to a new conversational QA scenario without direct exposure during training. As shown in Table 8, CompSelect consistently achieves strong performance in this unseen dataset, highlighting its robustness and ability to generalize across diverse QA tasks.
@@
[…]
6. Conclusion
In this work, we propose CompSelect, a compact clue selection mechanism for LLM-centric RAG that efficiently extracts, organizes, and truncates answer-relevant information from large-scale documents. By framing retrieval as a MinMax optimization, its three modules, the clue extractor, the reranker, and the truncator, work together to enhance reasoning quality and efficiency. Experiments on three QA datasets show that CompSelect improves performance and reduces latency, while remaining robust to unreliable retrieval and generalizing well across scenarios. Future work could explore integrating it with generative retrieval, allowing models to generate document indices, thereby avoiding costly online processing and further reducing end-to-end latency.
Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No.U25B2070 and No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
@@ A.1. Prompt for the Clue Extractor
We present our prompt for the clue extractor in Table 9. The prompt is designed to guide the model in extracting the most informative sentences, which are most likely to contain the answer to the given question.
@@ A.2. Prompt for the Adaptive Truncator
We present our prompt for the adaptive truncator in Table 10. The prompt is designed to guide the model in optimizing context truncation based on the complexity of the question and the quality of the retrieved documents, thereby improving the efficiency of the language model. Specifically, given a question and a ranked list of sentences, the model’s task is to identify and retain the most relevant sentences while truncating those that are irrelevant to the question.
@@ B.2. More Implementation Details
We use LLaMA3-8B-Instruct  [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Dubey et al., 2024), Qwen3-14B  [^5] https://huggingface.co/Qwen/Qwen3-14B (Yang et al., 2025), and LLaMA3.3-70B-Instruct  [^6] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct (Dubey et al., 2024) as the generators, covering medium, large, and ultra-large LLMs. All of these models demonstrate strong performance across various tasks and exhibit high flexibility. To reduce computational cost and latency, our clue extractor and adaptive truncator are based on LLaMA3.2-3B-Instruct  [^7] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct (Dubey et al., 2024). We apply the LoRA method (Hu et al., 2021) for fine-tuning, an efficient low-rank adaptation technique that reduces the computational cost of parameter updates while maintaining model performance. LoRA is applied to both the clue extractor and adaptive truncator. We train the models on two NVIDIA RTX PRO 6000 GPUs for 12 epochs. The initial learning rate is set to $1\times 10^{-4}$, and the batch size is 4. Gradient accumulation is employed to simulate larger effective batch sizes and improve training stability. The best model is selected based on validation set performance. During fine-tuning, the models are trained on the three QA datasets using a KNN-based sentence selection method. Data preprocessing is accelerated with 16 parallel workers per training epoch. The maximum input length is set to 4096 tokens to accommodate long-context information. For the clue reranker, we employ Sentence-BERT (Reimers and Gurevych, 2020) with the distilbert-base-uncased model to generate high-quality sentence embeddings for computing sentence similarity. Training uses the Adam optimizer with a batch size of 64, a learning rate of $2\times 10^{-5}$, and 1000 warm-up steps, and runs for 4 epochs.